1. A volume of carbon dioxide gas, CO2, equal to 25.0 L was collected at 23 0C and 98.9 kPa.  What would be the volume of carbon dioxide if it were collected at 23 0C and 83.0 kPa?
1
Expert's answer
2021-04-23T07:31:24-0400

"P_1=98.9kPa"

"V_1=25.0L"

"T_1=23+273=296k"


"P_2=83.0kPa"

"T_2=23.0+273= 296k"

"V_2=?"


Note that temperature is constant

Hence suitable formulae is

"{P_1\u00d7V_1}={P_2\u00d7V_2}"


"98.9\u00d725=83.0\u00d7V_2"


"V_2=\frac{98.9×25}{83.0}=29.79L"